Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
5 - 3z2 = 6z
HELP PLEASE

Answer by edjones(8007)

The discriminant = b^2-4ac
5 - 3z2 = 6z
y=3z^2-6z-5
a=3, b=-6, c=-5
(-6)^2-4*3*-5
=36-(-60)
=36+60
=96
.
Since the discriminant is greater than zero there are 2 real answers.
Since it is not a perfect square then the answers will be irrational.
.
Ed